A storage tank has shape of a circular cylinder .Suppose that the radius and height of the tank are measured at 1.5 m and 5m respectively with a possible error in calculating the capacity of the tank


1
Expert's answer
2021-11-30T11:41:21-0500
"V=\pi r^2h"

"dV=2\pi r h dr+\pi r^2dh"

"\Delta V=\pi r (2 h\Delta r+ r\Delta h)"

"\Delta V=\pi(1.5)(2(5)(0.05)+(1.5)(0.1))"

"\approx 3.063({ft}^3)\approx3.1{ft}^3"